An electron with a speed of 7.76 x 108 cm/s in the positive direction of an x axis enters an electric field of magnitude 1.53 x 103 N/C, traveling along a field line in the direction that retards its motion. (a) How far will the electron travel in the field before stopping momentarily, and (b) how much time will have elapsed? (c) If the region containing the electric field is 7.18 mm long (too short for the electron to stop within it), what fraction of the electrons initial kinetic energy will be lost in that region? (a) NumbeT 1119241 (b) Number T[2.88E-6 (c) Numbe UnitsT m Unit UnitsT This answer has no units
Answer #1

acceleration of electron

a = qE / m = 1.6*10^-19* 1530 / ( 9.1*10^-31)

a = 2.69*10^14 m/s^2

a) correct

b) t = u / a = 7.76*10^6 / ( 2.69*10^14) = 2.88*10^-8 seconds

c)

final velocity of electron

v'^2 = v^2 - 2 a d

v'^2 = (7.76*10^6)^2 - 2* 2.69*10^14* 7.18*10^-3

v' = 7.507*10^6 m/s

% energy lost = 0.5* m ( v^2 - v'^2)/ ( 0.5 m v^2)

% energy lost = (7.76 - 7.507)/ 7.76 = 0.0326

% Energy lost = 3.26 %

=========

Do comment in case any doubt will reply for sure.. Goodluck